OpenStudy (anonymous):

hoose the correct description of the graph of the compound inequality x - 1 < or = to 9 or 2x > or = to 24 A number line with an open circle on 10, shading to the left, and an open circle on 12, shading to the right. A number line with an open circle on 10, an open circle on 12, and shading in between. A number line with a closed circle on 10, a closed circle on 12, and shading in between. A number line with a closed circle on 10, shading to the left, and a closed circle on 12, shading to the right

OpenStudy (anonymous):

@dmezzullo sorry for taggiin you soo much..i know you must hate me right now ):

OpenStudy (dmezzullo):

Its ok idk this one

OpenStudy (anonymous):

poo

OpenStudy (anonymous):

\[x-1\leq 9\iff x\leq 10\]

OpenStudy (anonymous):

huh????????????

OpenStudy (anonymous):

\[2x\geq 24\iff x\geq 12\]

OpenStudy (anonymous):

im confsed!!

OpenStudy (anonymous):

before you can answer the question, you have to solve each inequality for \(x\) that is the first step

OpenStudy (anonymous):

the first inequality is \(x-1\leq 9\) to solve this for \(x\) add 1 to both sides, which gives you \[x\leq 10\]

OpenStudy (anonymous):

to solve \(2x\geq 24\) divide both sides by 2 and get \(x\geq 12\)

OpenStudy (anonymous):

this tells you \(x\leq 10\) OR \(x\geq 12\)

OpenStudy (anonymous):



OpenStudy (anonymous):

both circles are closed because you have \(\leq\) and \(\geq\) instead of \(<\) and \(>\)